OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y <= x - 5 y => -x - 4

OpenStudy (lalaly):

the following points?

OpenStudy (anonymous):

(-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (lalaly):

second

OpenStudy (anonymous):

Explain how you got the answer

OpenStudy (lalaly):

first point (-5,2) is 2<= -5-5 ... 2<= -10 it isnt, so u try the nxt point (5,-2) is -2<=5-5 .... -2<=0? it is, so try it on the second inequality is -2 >=-5-4.... -2>= -9 ye it is so the second point is the right one